# The unified non-disciplinary model of the operation of neurons

### v3

@@ A.1 Science methods
The confusion and question marks in connection with describing life by science mostly arise from the interpretation of notion ’speed’ in physics. When discussing the underlying physical laws for biology, we go back to the fundamental physical concepts instead of taking over the approximations and abstractions used in the classical physics for non-biological matter and less complex interactions. As we emphasized many times, we construct laws and conclusions based on simplified abstractions about nature in all fields of science. The notions and laws depend on the circle of phenomena we know and want to describe. The Newtonian and Einsteinian worlds are basically distinguished by considering speed dependence that actually means explicit time dependence. Interesting consequences are that in the Einsteinian world, the mass is not constant, time and space are not absolute, and so on. We can be prepared for some similar counter-intuitive experiences in physiology: "we must be prepared to find it working in a manner that cannot be reduced to the ordinary laws of physics"[7]. Here, we scrutinize the fundamental concepts and discover differences between physics and biology as consequences of the required different abstractions and approximations.
During our college studies, we mentioned that light is an electromagnetic wave with a vast but finite propagation speed. Still, we forgot to highlight that, at the same time, it is the propagation speed of the electric (, and magnetic, and gravitational) interaction force fields as well. The effect of "Retarded-Time Potential" is also known in physics and communication engineering. Algorithms "marching-on-in time" and "Analytical Retarded-Time Potential Expressions" are derived to handle the problem; for a discussion, see [88]. Telegrapher’s equations (unfortunately, also used to describe biological signal transfer) explicitly assume a finite propagation speed, millions of times slower than the (implicitly assumed) EM interaction’s. The issue is not confined to large distances: designers of micro-electronic devices also must consider the effect: they introduced clock time domains and clock distribution trees; see, for example [89, 90].
@@
Science uses ’instant’ in the sense that one interaction is much faster than the process under study; we consider the faster interaction as instant. The approach of classical science is based on the oversimplified approximation that the interaction speed is always much higher than the speed of changes it causes and that the processes can always be described by a single stage. In our approach, for biology, we put together a series of stages to describe the observed complex phenomena, where the stages provide input and output for each other, involve more than one interaction speed, and use per-stage-valid approximations. We simplify the approximations by omitting the less significant interactions and introduce ideas for accounting for the different interaction speeds. This way, we reduce the problem to a case that science can describe mathematically. This procedure is fundamentally different from applying some mathematical equations derived for an abstracted case of science to a complex biological phenomenon without validating that we use the appropriate formalism.
@@ A.2 Complete measurement
In physics, measurement means the determination of something relevant to the process under study. A ’complete’ measurement measures all relevant quantities. The different disciplines of physics restrict the measured quantities to the ones which are, in general, that the discipline studies. The remaining quantities remain outside the scope of the discipline. The fundamental physical quantities of mechanics are length, mass, and time, which form the basis for defining all other quantities like velocity, force, and energy in that field. The fundamental physical quantities in thermodynamics are Temperature, Energy (Internal Energy, Heat, Work), and Entropy, which characterize systems at equilibrium and describe energy transformations; Pressure and Volume also being key variables. Those of electricity include electric charge (Coulomb), the basis of all electric phenomena; electric current (Ampere), the rate of charge flow; voltage (Volt), the potential difference driving current; resistance (Ohm), opposition to current; power (Watt), the rate of energy transfer; and energy (Joule), the capacity to do work. As seen, there is little overlap of the studied quantities. A disciplinary study (such as the electrical and thermodynamic ones in physiology) provides an incomplete set of measured quantities. None of the disciplines alone can describe electrolytes given that they consider a different and incomplete set of physical quantities. As discussed in connection with the Onsager relations, a disciplinary (incomplete) measurement does not discover that an unexpected change (a miracle) in the value of one of the quantities belonging to another discipline is accompanied by a change in the value of the studied quantity. This change is clearly the case with the measurements performed in the spirit of HH: the values the electrical instruments provide are accompanied by the values of mechanical/thermodynamic quantities which are not measured, partly due to the obvious measuring difficulties and because they are outside the scope of the discipline. One cannot reduce ion-related phenomena to a single abstraction as thermodynamics does with mass or electricity does with charge. If one attempts (see the mentioned theoretical descriptions), one experiences that some quantity (the charge or the mass) changes in a uncontrolled way. As Feynman R. P [91] told, "the separation of fields, as we have emphasized, is merely a human convenience, and an unnatural thing. Nature is not interested in our separations, and many of the interesting phenomena bridge the gaps between fields". The disciplinary separation in classical science does not apply to the study of life.
@@ A.3 Physics and biology
Science could serve as a firm base for all its disciplines. As we discuss, its disciplines use abstractions based on limited-validity approximations based on the same first principles. However, the approximations differ between biology and physics. In physics, some processes we observe are fast enough so that we can use the abstraction that they are essentially state transitions. In some cases, the approach can be –more or less– successful. For the slower, well-observable processes, we have the laws of motion that describe how processes occur under the influence of a driving force. We also found that nature is not necessarily linear (in the sense that it depends only on the mentioned quantities, not on their derivatives), which we can describe with "nice" mathematical formulas. A century ago, A. Einstein discovered that the approximations I. Newton introduced two centuries earlier are not sufficiently accurate for describing the movement of bodies at high speeds. In other words, a new paradigm, the constancy of the speed of light, must have been introduced that caused a revolution in physics and led to the birth of "modern physics".
@@
Another point is that science started with the assumption that the non-living matter is continuous, although it was early discovered that there are the smallest pieces of that matter. When we reached that size, we experienced that different subsets of science laws describe that matter and the atoms they contain. It is one of the most challenging tasks to establish relations between those subsets. Again, we used abstractions that the continuous matter is infinitely large and that the isolated atoms are infinitely far from each other and from the external world. We also experienced the semi-infinite cases, and studied the behavior of surfaces and interfaces, which, again, is different from both that of the atoms and their large masses. Given that biological objects span the microscopic to macroscopic size range and are surrounded by surfaces, we must be prepared for the fact that no simple rules describe their behavior.
@@
Appendix B Speculations
$F_{Na^{+}}=10^{7}*1.60217663*10^{-19}[V][C]=1.6*10^{-12}\ [N]$ (52)
Again, the pressure can be calculated using electricity.
[…]
A way to estimate the ions’ rush-in speed is that one assumes that the Stokes-Einstein speed describes the ions’ speed (Fig. 3 in [14], quantitatively underpins the hypothesis, see [94]) both in axons and ion channels. If so, the ratio of the potential gradients equals the ratio of the speeds of ions. One can estimate that the 100mV100~mV AP on the $50~\mu m$ AIS generates a $2*10^{4}~\frac{V}{m}$ electrical gradient and [14] measured $20~\frac{m}{s}$ for the speed of the AP. By assuming that the electrical field across the membrane is $10^{7}~\frac{V}{m}$, one can estimate the speed at the exit of the ion channels as $10^{4}~\frac{m}{s}$.
We can also estimate the work done when the rush-in charge extends the neuron, as the $\Delta P\times\Delta V$. We consider that $\Delta P$ remains constant, and we calculate the change of the volume as neuron’s surface $10^{-8}\ m^{2}$ multiplied by the change of the radius $\Delta r=10^{-9}\ [m]$ multiplied by the pressure $\Delta P$, that gives $E_{mech}=2*10^{-14}\ [J]$. If one assumes that the pressure is of entirely mechanical origin [80], one must assume that the pressure is $10^{6}~\frac{N}{m^{2}}$, which is well above the measured value. We can also estimate the electrical energy $E_{electr}=\frac{1}{2}C*(\Delta V)^{2}=1.4*10^{-12}\ [J]$, see also [45]. To estimate the time required to change the pressure, one can assume that the collision force of a $Na^{+}$ ion provides the pressure (the shock wave), and this can be calculated as $F=m\frac{\Delta v}{\Delta t}$. Given that Eq. (52) provides the force, and one must assume that the ion loses its $10^{4}~to~10^{5}\frac{m}{s}$ speed, one concludes that it generates a shock wave in $\Delta t=10^{-8}~to~10^{-7}~seconds$.
In line with the experimentally measured energy consumption and our calculation in section 7, the AP cycle consumes energy in the range of $10^{-7}\ [J]$. The five orders of magnitude between the total and the electrical energy seems to underpin that in addition to the $10^{-3}$ concentration of ions, only a small fraction of the ions (the ones in the layers near the membrane) in the volume participate in the electric activity of issuing an AP, while the entire volume in the mechanical activity. That is, practically all the energy is invested as elastic energy, but the elasticity modulus of the membrane is exceptionally high. The primary reason for neuronal operation is the appearance of new ions in the vicinity of the membrane, but this creates a secondary, vast mechanical shock wave. By considering the electrical repulsion force between particles in a slow current, which is entirely omitted in the classic theory, we open the way for explaining the observed thermodynamic and mechanical changes. Instead of alternative disciplinary theories, cooperation between the classic disciplines is needed.
For thermodynamic distributions, one can interpret the "temperature of individual particles" with mean kinetic energy $E$ as $T=\frac{2*E}{3*k_{B}}$ (where $k_{B}$ is the Boltzmann constant and $T$ is the thermodynamic temperature of the bulk quantity of the substance); that is, the temperature is directly proportional to the average kinetic energy. The energy arrives at the neuron in the form of $10^{7}$ $Na^{+}$ ions, so the average energy a single ion consumes is $2.5*10^{-14}~[J]$. One can assume that most (99%) of that energy is consumed for moving the ion against the Stokes-Einstein force in the viscous medium, so after exiting the membrane, the speed of the ions is $\sqrt{\frac{2*2.5*10{-16}}{3.82*10^{-26}}}\approx 10^{5}~[m/s]$, which is well above the $\approx~500~m/s$ average speed. The ions on the two sides of the membrane are in the state with temperature $T$ before and after the AP. When the $Na^{+}$ ions rush into the intracellular space, they gain energy through electrostatic acceleration by the membrane’s electrical potential, so the same ions appear on the intracellular side as having more energy, i.e., slightly higher temperature (the temperature is more than two orders of magnitude higher, but the proportion of those "hot" ions is about six orders of magnitude lower, so one can expect temperature change up to the range of fragments of up to millikelvin, depending on the measuring conditions). In the second phase of the AP, those more energetic ions that provide excess voltage (and so: excess local potential) above the resting potential leave the membrane’s proximity through the AIS, so the temperature decreases in the second phase to its original value; another effect that can contribute to the general heat production and adsorption discussed above. In the original publication on solitons (essentially pressure waves) [15], a wide range of measured and theoretical data is discussed. They are in line with our conclusions, among others, $80\ \mu K$ measured temperature change. Furthermore, the electrical energy is more than an order of magnitude smaller than the elastic energy (called soliton energy there).

### v4

@@ A.1 Science methods
The confusion and question marks in connection with describing life by science mostly arise from the interpretation of notion ’speed’ in physics. When discussing the underlying physical laws for biology, we go back to the fundamental physical concepts instead of taking over the approximations and abstractions used in the classical physics for non-biological matter and less complex interactions. As we emphasized many times, we construct laws and conclusions based on simplified abstractions about nature in all fields of science. The notions and laws depend on the circle of phenomena we know and want to describe. The Newtonian and Einsteinian worlds are basically distinguished by considering speed dependence that actually means explicit time dependence. Interesting consequences are that in the Einsteinian world, the mass is not constant, time and space are not absolute, and so on. We can be prepared for some similar counter-intuitive experiences in physiology: "we must be prepared to find it working in a manner that cannot be reduced to the ordinary laws of physics"[9].
@@
Science uses ’instant’ in the sense that one interaction is much faster than the process under study; we consider the faster interaction as instant. The approach of classical science is based on the oversimplified approximation that the interaction speed is always much higher than the speed of changes it causes and that the processes can always be described by a single stage. In our approach, for biology, we put together a series of stages to describe the observed complex phenomena, where the stages provide input and output for each other, involve more than one interaction speed, and use per-stage validity. We simplify the approximations by omitting the less significant interactions and introduce ideas for accounting for the different interaction speeds. This way, we reduce the problem to a case that science can describe mathematically. This procedure is fundamentally different from applying some mathematical equations derived for an abstracted case of inanimate science to a complex biological phenomenon without validating that we use the appropriate formalism.
@@ A.2 Complete measurement
In physics, measurement means the quantization of something relevant to the process under study. A ’complete’ measurement measures all relevant quantities. The different disciplines of physics restrict the measured quantities to the ones which are, in general, that the discipline studies. The remaining quantities remain outside the scope of the discipline. The fundamental physical quantities of mechanics are length, mass, and time, which form the basis for defining all other quantities like velocity, force, and energy in that field. The fundamental physical quantities in thermodynamics are Temperature, Energy (Internal Energy, Heat, Work), and Entropy, which characterize systems at equilibrium and describe energy transformations; Pressure and Volume are also key variables. Those of electricity include electric charge (Coulomb), the basis of all electric phenomena; electric current (Ampere), the rate of charge flow; voltage (Volt), the potential difference driving current; resistance (Ohm), opposition to current; power (Watt), the rate of energy transfer; and energy (Joule), the capacity to do work. As seen, there is little overlap of the studied quantities. A disciplinary study (such as the electrical and thermodynamic ones in physiology) provides an incomplete set of measured quantities. None of the disciplines alone can describe electrolytes, since they consider different and incomplete sets of physical quantities. As discussed in connection with the Onsager relations, a disciplinary (incomplete) measurement does not discover that an unexpected change (a miracle) in the value of one of the quantities belonging to another discipline is accompanied by a change in the value of the studied quantity. This change is clearly the case with the measurements performed in the spirit of HH: the values the electrical instruments provide are accompanied by the values of mechanical/thermodynamic quantities which are not measured, partly due to the obvious measuring difficulties, and because they are outside the scope of the discipline. One cannot reduce ion-related phenomena to a single abstraction as thermodynamics does with mass or electricity does with charge. If one attempts (see the mentioned theoretical descriptions), one experiences that some quantity (the charge or the mass) changes in an uncontrolled way.
@@ A.3 Physics and biology
Science could serve as a firm base for all its disciplines. As we discuss, its disciplines use abstractions based on limited-validity approximations based on the same first principles. However, the approximations differ between biology and physics. In physics, some processes we observe are fast enough that we can treat them as essentially state transitions. In some cases, the approach can be –more or less– successful. For the slower, well-observable processes, we have the laws of motion that describe how processes occur under the influence of a driving force. We also found that nature is not necessarily linear (in the sense that it depends only on the mentioned quantities, not on their derivatives), which we can describe with "nice" mathematical formulas. A century ago, A. Einstein discovered that the approximations I. Newton introduced two centuries earlier are not sufficiently accurate for describing the movement of bodies at high speeds. In other words, a new paradigm, the constancy of the speed of light, must have been introduced that caused a revolution in physics and led to the birth of "modern physics".
@@
Another point is that science started with the assumption that the non-living matter is continuous, although it was early discovered that there are the smallest pieces of that matter. When we reached that size, we experienced that different subsets of science laws describe that matter and the atoms they contain. Establishing relations between those subsets is one of the most challenging tasks. Again, we used abstractions that the continuous matter is infinitely large and that the isolated atoms are infinitely far from each other and from the external world. We also experienced the semi-infinite cases, and studied the behavior of surfaces and interfaces, which, again, is different from both that of the atoms and their large masses. Given that biological objects span the microscopic to macroscopic size range and are surrounded by surfaces, we must be prepared for the fact that no simple rules describe their behavior.